# Information Discovery in e-Commerce

### v1

@@ 4.2 User profiling in e-commerce
A user profile refers to personal information about a specific user. Personalization plays an important role in web search and recommender systems. In e-commerce portals, user profiling is a critical module for e-commerce information discovery tasks, as it provides personalized content in search or recommendation results. User profiling can be defined as the process of exploring information about a user’s interest domain [dong2014inferring, kanoje2015user]. Information about a user can be used by e-commerce search and recommender systems to enhance the system effectiveness because it enables better user understanding (see Section 5.3.5 and  6.4). Given that it originated from work on the prediction of user purchase intention, research into user profiling in e-commerce has continuously garnered attention over the years [see, e.g., solomon1994buying, braynov2003personalization, hollerit2013towards, zhang2013predicting, gupta2014identifying, rahdari2017analysis, Huangwww2018]. Early work on user profiling for e-commerce mainly focused on information filtering, social media analysis, web searches, and fraud detection [solomon1994buying, fawcett1996combining, adomavicius1999user, kuflik2000generation, braynov2003personalization]. Most of these are rule-based strategies. For example, fawcett1996combining employed a rule-based user-profiling method to uncover indicators of fraudulent behavior. These indicators were used to create user profiles that were then applied as features of their proposed system, which combines evidence from multiple profilers to generate high-confidence alarms.
@@ 4.2.1 Types of user profiling
User profiling can be classified as either profile extraction and profile learning [tang2010combination]. Profile extraction focuses on extracting information about a user, such as demographic data (e.g., age, gender, location) or basic behavior patterns. However, in the environment of e-commerce, profile extraction is often less critical because it provides only a fixed snapshot of the user, lacking the adaptability needed to capture evolving preferences and real-time behavioral changes. In contrast, profile learning is more significant as an e-commerce modeling tool [kuflik2000generation, cufoglu2014user]. Profile learning methods can be grouped into three categories: (i) content-based methods, (ii) collaborative methods, and (iii) hybrid methods [cufoglu2014user]. Content-based methods infer user profiles based solely on the users’ own previous behavior [kuflik2000generation]. Collaborative methods in user profiling focus on applying collaborative filtering approaches to infer profile information based on the behavior of users in a suitably defined neighborhood of similar users. Collaborative filtering methods can be classified as either memory-based methods or model-based methods [godoy2005user, cufoglu2014user]. Memory-based solutions estimate ratings for a user based on the entire collection of previous ratings of similar users [adomavicius2005toward, su2009survey]. In contrast to memory-based collaborative filtering methods for user profiling, model-based methods use the collection of ratings to learn a model to estimate user profiles [su2009survey]. Hybrid methods have garnered attention because they combine content-based methods and collaborative methods [godoy2005user]. Regarding research on personalized recommendations, to capture users’ information and interest, liu2010iui proposed a dynamic collaborative filtering method for news recommendation where the recommender constructs user profiles based on their past click behavior. The authors initially conducted a log analysis of the changes in user interest in news topics over time. By classifying users’ news interests as either genuine interests or the influence of local news trends, the authors were able to (i) construct a Bayesian framework to model a user’s genuine interests based on their past click behavior, and (ii) predict current interests by jointly analyzing the genuine interest and the local news trends.
@@ 4.2.2 User profiling with social media
Social media is playing an important role in e-commerce user profiling. On the one hand, social media provides a source for generating user profiles, especially when addressing the cold start problem. mishne-deriving-2006 provided an early example of this idea, using a combination of text analysis and external knowledge sources to estimate the commercial tastes of bloggers from their posts. On the other hand, profiling information learned from social media data can be applied to explore the user’s purchase intentions on e-commerce platforms. Targeting users with no history on an e-commerce site, zhang2013predicting focused on predicting the purchase behavior by proposing a feature-selection method to predict the product categories from which a user will buy. Representing user purchase intent as textual information that indicates a desire to purchase a product or service in the future, gupta2014identifying proposed a binary classification approach to identify the user purchase intention based on their social media posts. ding2015mining explored relationships between a user’s consumption habits and their social media data. They proposed a consumption intention mining model (CIMM) based on CNNs. lo2016understanding analyzed user activities in social media to build a time-varying model to predict user purchase intent. The authors analyzed Pinterest [^1] https://www.pinterest.com data to understand how the usage of an e-commerce platform relates to future user shopping behavior. They found that indicators of purchase intent tended to gradually build up over time and sharply increase 3 to 5 days before purchase. Multi-modal information has also been applied to infer user profiles. gelli2017personality focused on automatically discovering actionable images for users according to their personality. By applying their model to a large-scale dataset, the authors found a significant correlation between personality traits and affective visual concepts in the image content.
@@ 4.2.3 Graph-based user profiling
Most approaches to user profiling only use a single type of information. In e-commerce modeling, heterogeneous graphs are also being used to work with user profiles. Fig. 4.6 shows an example of a graph with heterogeneous information; particularly, three kinds of nodes were applied to represent three types of data, i.e., users, items, and attributes, respectively. chen2019semi focused on applying rich interactions among data instances, i.e., co-click and co-purchase behavior on e-commerce platforms, to enhance user-profiling performance in e-commerce models. Neighborhood features were found to provide useful information that helps to infer user profiles, e.g., users that have similar co-purchase behavior on e-commerce platforms are likely to be of a similar age. The authors proposed a heterogeneous graph attention network to infer user profiles within a multi-type data environment; the network was able to model the rich unsupervised information in a heterogeneous graph by encoding the graph structure and node features. gu2020hierarchical constructed a hierarchical profiling framework to model users’ real-time interests at different granularities. A pyramid recurrent neural network model for hierarchical user profiling was constructed based on users’ micro-behavior; it was subsequently applied to model the types and dwell times of behavior to enable an effective formulation of users’ real-time interests. These graph-based user-profiling methods have been applied to enhance the re-ranking results in e-commerce recommendation systems. More details are discussed in Section 6.4.
@@ 4.3.1 Graph learning for user behavior modeling
Graph neural networks use neural networks to represent graph information [scarselli2009graph, bruna2013spectral, battaglia2016interaction]. Graph convolutional networks extend CNNs to graph structured data; they have been shown to be effective on a range of graph classification tasks [bruna2013spectral, defferrard2016convolutional, zhao2021sigirwgcn, liu2022sigirgraph], and when applied for semi-supervised classification [hamilton2017inductive, kipf2016semi] and link prediction [ma2018dynamic]. Most of these models have been designed for static graphs. However, ma2018dynamic proposed a dynamic graph neural network model that can model dynamic information. xu2021transformer enhanced the capacity for learning complicated temporal dependency in the graph, by proposing a transformer-style relational reasoning network with a dynamic memory updating mechanism. Graph neural networks that have been successfully applied to recommender systems were proven to be able to learn item embeddings within a large-scale item relationship graph [ying2018graph] that describes users, items, and pairwise relations in e-commerce scenarios [ma2018dynamic, gaowsdm2022rec]. Graph learning can also be applied for post-click modeling and user purchase-intent modeling. To predict fine-grained post-click CVR, bao2020gmcm designed a model to represent user micro-behavior as a purchase-related micro-behavior graph. The authors applied a multi-task learning framework to construct a graph-based micro-behavior conversion model that can capture the correlation between different types of micro-behavior. The proposed multi-task learning and inverse propensity weighting modules mitigate the data sparsity- and sample selection bias-related problems. To prediction efficient and accurate CVR, wen2021hierarchy proposed a graph neural network to hierarchically model both micro- and macro-behaviors in a unified framework. It seems likely that graph neural networks will continue to facilitate new ways of modeling, gaining insights into, and predicting, e-commerce user behavior in search (Section 5.3.3) and recommendation scenarios (Section 6.2.2).
@@ 4.3.2 Dynamic user behavior modeling and profiling
Most previous studies on e-commerce user behavior modeling and user profiling were conducted under the assumption that a snapshot of user behavior is recorded on e-commerce portals. However, a user’s personal interests and behavior may change over time [koren2009collaborative, gao2013modeling, yin2014temporal, huang2015tencentrec, yin2015dynamic, jagerman2019people, wang2021clicks, huang2022clickrec]. Thus, modeling e-commerce users’ temporal behavior is important for e-commerce search and recommendation system development. Social networks and social media provide a rich source of information for temporal models of user behavior [mislove2010you, gao2013modeling, yin2014temporal, yin2015dynamic]. yin2014temporal designed a latent mixture model, which they named a temporal context-aware mixture model, to account for the intentions and preferences that drive user behavior. It models the topics related to users’ intrinsic interests, and the topics related to temporal context’ it then jointly analyzes the influences of the two factors to model user behavior in a unified way. To enable the dynamic learning of user profiles, cao2017you developed a model that considers multiple information sources and their relations. Similarly, liang2018dynamic-kdd proposed a streaming profiling algorithm that initially applies a user-expertise tracking model to track the changes in the dynamic expertise of users; it then utilizes a keyword diversification algorithm to produce top-$k$ diversified keywords that allow the users’ dynamic expertise to be profiled at a specific timestamp.
@@ 4.3.3 User modeling with insufficient data
As we have discussed in this chapter, a wide range of click models and ranking methods can be applied to model user click behavior [chuklin-click-2015, borisov2016neural, ferro2017including, he2017neural, wu2018turning, ferro2019boosting, liu2022sigirgraph, vardasbi-2022-probabilistic]. However, the models developed to date tend to require fully labeled data to train the ranking models, although, in realistic e-commerce scenarios, not all of a user’s behavior can be recorded. Similarly, regarding post-click modeling, purchase-intent prediction, and user-profiling tasks, the reality that there is a limited amount of behavioral data makes it difficult to work with existing click modeling and ranking solutions. For example, in the case of CVR tasks, the data sparsity problem arises when the number of training samples for the sequential behavior of the form “click $\to$ purchase” is insufficient to fit the large parameter space of the CVR task [wen2020entire, guo2021enhanced]. How to enhance user modeling performance under the conditions of a limited amount of imperfectly labeled data remains an important open problem in e-commerce.
@@
E-commerce search, or simply “product search,” represents a special retrieval scenario where users submit queries to retrieve products from a search engine [ai2017learning]. E-commerce search portals are gaining in popularity as many consumers choose e-commerce search on an e-commerce platform rather than generic web search [li2011towards]. Unlike in web search, in e-commerce search there can be millions of results to surface for a given search query [wu2018turning]. We have discussed user behavior modeling and user profiling in Chapter 4 to understand how to explore and exploit information from user behavior. In this chapter, we focus on the other side of the coin, on search technologies that are based on users’ interactive behavior. To learn about e-commerce search solutions, we discuss research on query understanding and ranking technologies for e-commerce search. In Section 5.1 we summarize characteristics of e-commerce search. In Section 5.2 we recall key metrics being used for evaluating e-commerce search. In Section 5.3 we present studies on representing e-commerce search queries. Then, we detail e-commerce ranking approaches in Section 5.4. Finally, we discuss emerging research directions in e-commerce search in Section 5.5.
@@ 5.1.1 Overview of e-commerce search
Early e-commerce search approaches are based on traditional information retrieval theory and faceted search models [yee2003faceted, jansen2006effectiveness]. jansen2006effectiveness explored the difference between ad-hoc search and e-commerce search. A grocery retrieval system was developed by considering a discrepancy between consumers’ shopping lists and retailers’ stock information [nurmi2008product]. Early on, e-commerce search systems relied on information that retailers make available: either semantic markup on unstructured HTML documents or a data feed provided in some predefined structured format. Product resolution [balog2011investigation] focuses on recognizing webpages that represent the same product. Based on the task of product resolution, duan2013probabilistic proposed a probabilistic mixture model for mining and analyzing product search logs. Similar setups can be also found in a product-aware keyword search system [duan2013supporting].
@@
Unlike traditional ad-hoc retrieval, e-commerce search relies on a decision mechanism about consumers’ purchase behavior in e-commerce portals [li2011towards]. There are two main stakeholders in e-commerce search, consumers and business owners, whose interests align but also conflict to a certain extent [tsagkias-2020-challenges]. In e-commerce search, customers do not just browse relevant items, but also try to locate an item that satisfies their specific purchase intent [li2011towards]. While consumers aim to find the best quality at the lowest price, businesses want to maximize profit, which translates into higher prices for customers or lower costs for businesses. E-commerce search typically requires more structured information (e.g., brands, categories, shops, etc.) than web search and more diversified personal definitions of “relevance” during search sessions. On the one hand, as we have discussed in Section 3.2, users come to e-commerce websites with a wide spectrum of intents. Hence, multiple user behavior discussed in Chapter 3 and Chapter 4, e.g., clicks, post-clicks, and purchases, etc., should be integrated to model the “relevance” in e-commerce search. On the other hand, only a few products are actually purchased by the consumers and different individuals have different opinions even about the same product [ai2017learning]. Thus, e-commerce search should consider users’ differences to satisfy the needs of all consumers. In general, there are four unique characteristics in e-commerce search:
@@
  • Consumer query intent. Similar to web search, queries in e-commerce search can be divided into three classes: navigational, informational and transactional [li2011towards]. However, e-commerce search queries take a different form. Specifically, navigational queries are product serial numbers and inquiries for customer support; informational queries include leaves in the product taxonomy and product attributes; and transactional queries are a mix of navigational and informational queries. Unlike traditional web search, there are three query intents for e-commerce search: target finding, decision making, and exploration [su2018user]. Following the well-known web-search taxonomy due to broder2002taxonomy, su2018user proposed a hierarchical e-commerce search taxonomy to explore consumers’ shopping intents, with shallow exploration, targeted purchase, major-item shopping, minor-item shopping, and hard-choice shopping. The authors found that consumers tend to conduct more focused searches in target finding sessions compared to those in the decision making and exploration sessions. In target finding sessions, consumers tend to issue a few specific queries and browse only top ranked results; in decision making sessions, consumers tend to issue short queries, browse deep, and click more results; and in exploration sessions consumers issue many diverse queries but do not click often. Given all these search intents, customized search approaches for each type of search queries can be developed to improve the utility of e-commerce search.
@@
  • Heterogeneous consumer behavior. As we have described in Chapter 3, multiple types of user behavior can be observed in e-commerce platforms. During an online shopping journey, a consumer may have multiple targets at different stages. During a journey, blake2016returns found that e-commerce search proceeds as a kind of “funnel” where, initially, search is along broad categories, and then it becomes more refined to obtain an item at the lowest cost given a consumer’s cost of search. Hence, e-commerce search approaches should be aware of the stages in each consumer’s journey. Meanwhile, the overall impact of heterogeneous consumer behavior also makes e-commerce search different from traditional web search. Users’ micro behavior, post-click behavior, and engagement make the search intent dynamic and complicated during search sessions [zhouwsdm2018, wu2018turning].
@@
  • Online and offline ranking. Traditional learning to rank methods sort documents according to their relevance to the query. E-commerce search has an intrinsic difference in the relevance in rankings: the notion of “relevance” is blurred [wu2018turning]. Users come to e-commerce platforms with a wide spectrum of intents. Some users wish to make a purchase as soon as possible while others are just wandering around the platform to get inspired. Hence, various kinds of signals, including clicks, favorites, adding carts, purchases, etc., should be integrated to model the relevance in e-commerce search. E-commerce businesses having both online and physical presence bring a unique blend of infrastructure challenges [ariannezhad-2021-understanding]. Thus, users’ shopping experiences lie in smooth transitions from offline to online and vice versa [tsagkias-2020-challenges].

### v3

@@ 4.2 User profiling in e-commerce
A user profile refers to personal information about a specific user. Personalization plays an important role in web search and recommender systems. In e-commerce portals, user profiling is a critical module for e-commerce information discovery tasks, as it provides personalized content in search or recommendation results. User profiling can be defined as the process of exploring information about a user’s interest domain [dong2014inferring, kanoje2015user]. Information about a user can be used by e-commerce search and recommender systems to enhance the system effectiveness because it enables better user understanding (see Section 5.3.5 and 6.4). Given that it originated from work on the prediction of user purchase intention, research into user profiling in e-commerce has continuously garnered attention over the years [see, e.g., solomon1994buying, braynov2003personalization, hollerit2013towards, zhang2013predicting, gupta2014identifying, rahdari2017analysis, Huangwww2018]. Early work on user profiling for e-commerce mainly focuses on information filtering, social media analysis, web searches, and fraud detection [solomon1994buying, fawcett1996combining, adomavicius1999user, kuflik2000generation, braynov2003personalization]. Most of these are rule-based strategies. For example, fawcett1996combining employ a rule-based user-profiling method to uncover indicators of fraudulent behavior. These indicators were used to create user profiles that were then applied as features of their proposed system, which combines evidence from multiple profilers to generate high-confidence alarms.
@@ 4.2.1 Types of user profiling
User profiling can be classified as either profile extraction and profile learning [tang2010combination]. Profile extraction focuses on extracting information about a user, such as demographic data (e.g., age, gender, location) or basic behavior patterns. However, in e-commerce, profile extraction is often less critical because it provides only a fixed snapshot of the user, lacking the adaptability needed to capture evolving preferences and real-time behavioral changes. In contrast, profile learning is more significant as an e-commerce modeling tool [kuflik2000generation, cufoglu2014user]. Profile learning methods can be grouped into three categories: (i) content-based methods, (ii) collaborative methods, and (iii) hybrid methods [cufoglu2014user]. Content-based methods infer user profiles based solely on the users’ own previous behavior [kuflik2000generation]. Collaborative methods in user profiling focus on applying collaborative filtering approaches to infer profile information based on the behavior of users in a suitably defined neighborhood of similar users. Collaborative filtering methods can be classified as either memory-based methods or model-based methods [godoy2005user, cufoglu2014user]. Memory-based solutions estimate ratings for a user based on the entire collection of previous ratings of similar users [adomavicius2005toward, su2009survey]. In contrast to memory-based collaborative filtering methods for user profiling, model-based methods use the collection of ratings to learn a model to estimate user profiles [su2009survey]. Hybrid methods have garnered attention because they combine content-based methods and collaborative methods [godoy2005user]. Regarding research on personalized recommendations, to capture users’ information and interest, liu2010iui introduce a dynamic collaborative filtering method for news recommendation where the recommender constructs user profiles based on their past click behavior. The authors conduct a log analysis of the changes in user interest in news topics over time. By classifying users’ news interests as either genuine interests or the influence of local news trends, the authors are able to (i) construct a Bayesian framework to model a user’s genuine interests based on their past click behavior, and (ii) predict current interests by jointly analyzing the genuine interest and the local news trends.
@@ 4.2.2 User profiling with social media
Social media is playing an important role in e-commerce user profiling. On the one hand, social media provides a source for generating user profiles, especially when addressing the cold start problem. mishne-deriving-2006 provided an early example of this idea, using a combination of text analysis and external knowledge sources to estimate the commercial tastes of bloggers from their posts. On the other hand, profiling information learned from social media data can be applied to explore the user’s purchase intentions on e-commerce platforms. Targeting users with no history on an e-commerce site, zhang2013predicting focus on predicting the purchase behavior by proposing a feature-selection method to predict the product categories from which a user will buy. Representing user purchase intent as textual information that indicates a desire to purchase a product or service in the future, gupta2014identifying propose a binary classification approach to identify the user purchase intention based on their social media posts. ding2015mining explore relationships between a user’s consumption habits and their social media data. They propose a consumption intention mining model (CIMM) based on CNNs. lo2016understanding analyze user activities in social media to build a time-varying model to predict user purchase intent. The authors analyze Pinterest [^1] https://www.pinterest.com data to understand how the usage of an e-commerce platform relates to future user shopping behavior. They find that indicators of purchase intent tended to gradually build up over time and sharply increase 3 to 5 days before purchase. Multi-modal information has also been applied to infer user profiles. gelli2017personality focus on automatically discovering actionable images for users according to their personality. By applying their model to a large-scale dataset, the authors find a significant correlation between personality traits and affective visual concepts in the image content.
@@ 4.2.3 Graph-based user profiling
Most approaches to user profiling only use a single type of information. In e-commerce modeling, heterogeneous graphs are also being used to work with user profiles. Fig. 4.6 shows an example of a graph with heterogeneous information; particularly, three kinds of nodes were applied to represent three types of data, i.e., users, items, and attributes, respectively. chen2019semi focus on applying rich interactions among data instances, i.e., co-click and co-purchase behavior on e-commerce platforms, to enhance user-profiling performance in e-commerce models. Neighborhood features provide useful information that helps to infer user profiles, e.g., users that have similar co-purchase behavior on e-commerce platforms are likely to be of a similar age. The authors propose a heterogeneous graph attention network to infer user profiles within a multi-type data environment; the network is able to model unsupervised information in a heterogeneous graph by encoding the graph structure and node features. gu2020hierarchical construct a hierarchical profiling framework to model users’ real-time interests at different granularities. A pyramid recurrent neural network model for hierarchical user profiling is constructed based on users’ micro-behavior; it is subsequently applied to model the types and dwell times of behavior to enable an effective formulation of users’ real-time interests. These graph-based user-profiling methods have been applied to enhance the re-ranking results in e-commerce recommendation systems. More details are discussed in Section 6.4.
@@ 4.3.1 Graph learning for user behavior modeling
Graph neural networks use neural networks to represent graph information [scarselli2009graph, bruna2013spectral, battaglia2016interaction]. Graph convolutional networks extend CNNs to graph structured data; they have been shown to be effective on a range of graph classification tasks [bruna2013spectral, defferrard2016convolutional, zhao2021sigirwgcn, liu2022sigirgraph], and when applied for semi-supervised classification [hamilton2017inductive, kipf2016semi] and link prediction [ma2018dynamic]. Most of these models have been designed for static graphs. However, ma2018dynamic propose a dynamic graph neural network model that can model dynamic information. xu2021transformer enhance the capacity for learning complicated temporal dependencies in a graph, by proposing a transformer-style relational reasoning network with a dynamic memory updating mechanism. Graph neural networks that applied to recommender systems learn item embeddings within a large-scale item relationship graph [ying2018graph] that describes users, items, and pairwise relations in e-commerce scenarios [ma2018dynamic, gaowsdm2022rec]. Graph learning can also be applied for post-click modeling and user purchase-intent modeling. To predict fine-grained post-click CVR, bao2020gmcm design a model to represent user micro-behavior as a purchase-related micro-behavior graph. The authors apply a multi-task learning framework to construct a graph-based micro-behavior conversion model that can capture the correlation between different types of micro-behavior. The proposed multi-task learning and inverse propensity weighting modules mitigate the data sparsity- and sample selection bias-related problems. To prediction efficient and accurate CVR, wen2021hierarchy propose a graph neural network to hierarchically model both micro- and macro-behaviors in a unified framework. It seems likely that graph neural networks will continue to facilitate new ways of modeling, gaining insights into, and predicting, e-commerce user behavior in search (Section 5.3.3) and recommendation scenarios (Section 6.2.2).
@@ 4.3.2 Dynamic user behavior modeling and profiling
Most studies on e-commerce user behavior modeling and user profiling are conducted under the assumption that a snapshot of user behavior is recorded on e-commerce portals. However, a user’s personal interests and behavior may change over time [koren2009collaborative, gao2013modeling, yin2014temporal, huang2015tencentrec, yin2015dynamic, jagerman2019people, wang2021clicks, huang2022clickrec]. Thus, modeling e-commerce users’ temporal behavior is important for e-commerce search and recommendation system development. Social networks and social media provide a rich source of information for temporal models of user behavior [mislove2010you, gao2013modeling, yin2014temporal, yin2015dynamic]. yin2014temporal design a latent mixture model, which they named a temporal context-aware mixture model, to account for the intentions and preferences that drive user behavior. It models the topics related to users’ intrinsic interests, and the topics related to temporal context’ it then jointly analyzes the influences of the two factors to model user behavior in a unified way. To enable the dynamic learning of user profiles, cao2017you developed a model that considers multiple information sources and their relations. Similarly, liang2018dynamic-kdd proposes a streaming profiling algorithm that initially applies a user-expertise tracking model to track the changes in the dynamic expertise of users; it then uses a keyword diversification algorithm to produce top-$k$ diversified keywords that allow the users’ dynamic expertise to be profiled at a specific timestamp.
@@ 4.3.3 User modeling with insufficient data
As we have discussed in this section, a wide range of click models and ranking methods can be applied to model user click behavior [chuklin-click-2015, borisov2016neural, ferro2017including, he2017neural, wu2018turning, ferro2019boosting, liu2022sigirgraph, vardasbi-2022-probabilistic]. However, the models developed to date tend to require fully labeled data to train the ranking models, although, in realistic e-commerce scenarios, not all of a user’s behavior can be recorded. Similarly, regarding post-click modeling, purchase-intent prediction, and user-profiling tasks, the reality that there is a limited amount of behavioral data makes it difficult to work with existing click modeling and ranking solutions. For example, in the case of CVR tasks, the data sparsity problem arises when the number of training samples for the sequential behavior of the form “click $\to$ purchase” is insufficient to fit the large parameter space of the CVR task [wen2020entire, guo2021enhanced]. How to enhance user modeling performance under the conditions of a limited amount of imperfectly labeled data remains an important open problem in e-commerce.
@@
E-commerce search, or simply “product search,” represents a special retrieval scenario where users submit queries to retrieve products using a search engine [ai2017learning]. E-commerce search portals are gaining in popularity as many consumers choose e-commerce search on an e-commerce platform rather than generic web search [li2011towards]. Unlike in web search, in e-commerce search there can be millions of results to surface for a given search query [wu2018turning]. We have discussed user behavior modeling and user profiling in Section 4 to understand how to explore and exploit information from user behavior. In this section, we focus on the other side of the coin, on search technologies that are based on users’ interactive behavior. To learn about e-commerce search solutions, we discuss research on query understanding and ranking technologies for e-commerce search. In Section 5.1 we summarize characteristics of e-commerce search. In Section 5.2 we recall key metrics used for evaluating e-commerce search. In Section 5.3 we present studies on representing e-commerce search queries. Then, we detail e-commerce ranking approaches in Section 5.4. Finally, we discuss emerging research directions in e-commerce search in Section 5.5.
@@ 5.1.1 Overview of e-commerce search
Early e-commerce search approaches are based on traditional information retrieval theory and faceted search models [yee2003faceted, jansen2006effectiveness]. jansen2006effectiveness explore the difference between ad-hoc search and e-commerce search. A grocery retrieval system has been developed by considering a discrepancy between consumers’ shopping lists and retailers’ stock information [nurmi2008product]. Early e-commerce search systems rely on information that retailers make available: either semantic markup on unstructured HTML documents or a data feed provided in some predefined structured format. Product resolution [balog2011investigation] focuses on recognizing webpages that represent the same product. Based on the task of product resolution, duan2013probabilistic propose a probabilistic mixture model for mining and analyzing product search logs. Similar setups can be also found in a product-aware keyword search system [duan2013supporting].
@@
Unlike traditional ad-hoc retrieval, e-commerce search relies on a decision mechanism about consumers’ purchase behavior in e-commerce portals [li2011towards]. There are two main stakeholders in e-commerce search, consumers and business owners, whose interests align but also conflict to a certain extent [tsagkias-2020-challenges]. In e-commerce search, customers do not just browse relevant items, but also try to locate an item that satisfies their specific purchase intent [li2011towards]. While consumers aim to find the best quality at the lowest price, businesses want to maximize profit, which translates into higher prices for customers or lower costs for businesses. E-commerce search typically requires more structured information (e.g., brands, categories, shops, etc.) than web search and more diversified personal definitions of “relevance” during search sessions. On the one hand, as we have discussed in Section 3.2, users come to e-commerce websites with a wide spectrum of intents. Hence, multiple user behavior discussed in Section 3 and 4, e.g., clicks, post-clicks, and purchases, etc., should be integrated to model the “relevance” in e-commerce search. On the other hand, only a few products are actually purchased by the consumers and different individuals have different opinions even about the same product [ai2017learning]. Thus, e-commerce search should consider users’ differences to satisfy the needs of all consumers. In general, there are four unique characteristics in e-commerce search:
@@
  • Consumer query intent. Similar to web search, queries in e-commerce search can be divided into three classes: navigational, informational and transactional [li2011towards]. However, e-commerce search queries take a different form. Specifically, navigational queries are product serial numbers and inquiries for customer support; informational queries include leaves in the product taxonomy and product attributes; and transactional queries are a mix of navigational and informational queries. Unlike traditional web search, there are three query intents for e-commerce search: target finding, decision making, and exploration [su2018user]. Following the well-known web-search taxonomy due to broder2002taxonomy, su2018user describe a hierarchical e-commerce search taxonomy to explore consumers’ shopping intents, with shallow exploration, targeted purchase, major-item shopping, minor-item shopping, and hard-choice shopping. The authors find that consumers tend to conduct more focused searches in target finding sessions compared to those in the decision making and exploration sessions. In target finding sessions, consumers tend to issue a few specific queries and browse only top ranked results; in decision making sessions, consumers tend to issue short queries, browse deep, and click more results; and in exploration sessions consumers issue many diverse queries but do not click often. Given these search intents, customized search approaches for each type of search queries can be developed to improve the utility of e-commerce search.
@@
  • Heterogeneous consumer behavior. As we have described in Section 3, multiple types of user behavior can be observed in e-commerce platforms. During an online shopping journey, a consumer may have multiple targets at different stages. blake2016returns observe that, during a journey, e-commerce search proceeds as a kind of “funnel” where, initially, search is along broad categories, and then it becomes more refined to obtain an item at the lowest cost given a consumer’s cost of search. Hence, e-commerce search approaches should be aware of the stages in each consumer’s journey. Meanwhile, the overall impact of heterogeneous consumer behavior also makes e-commerce search different from traditional web search. Users’ micro behavior, post-click behavior, and engagement make the search intent dynamic and complicated during search sessions [zhouwsdm2018, wu2018turning].
@@
  • Online and offline ranking. Traditional learning to rank methods sort documents according to their relevance to the query. E-commerce search has an intrinsic difference in the relevance in rankings: the notion of “relevance” is blurred [wu2018turning]. Users come to e-commerce platforms with a wide spectrum of intents. Some users wish to make a purchase as soon as possible while others are just wandering around the platform to get inspired. Hence, various kinds of signals, including clicks, favorites, adding carts, purchases, etc., should be integrated to model relevance in e-commerce search. E-commerce businesses having both online and physical presence bring a unique blend of infrastructure challenges [ariannezhad-2021-understanding]. Thus, users’ shopping experiences lie in smooth transitions from offline to online and vice versa [tsagkias-2020-challenges].